We are given 7 squares of side lengths 1, 2, 2, 2, 3, 4, 5. They can be fitted to form a rectangle, with no overlap or gaps in-between. What is the perimeter of the rectangle?

Grade:11

3 Answers

Akash Kumar Dutta
10 years ago

Dear Vivek,
Area of all the squares=area of rectangle=1.1 +3.4 +1.9+1.16+1.25=63
sice area is also l.b...
hence sides of rect.= 7 and 9.....[the sides cant be 1 and 63]
perimeter=7+9=16
ANS
Regards.

Akash Kumar Dutta
10 years ago

Dear Vivek,
a liitle mistake...

the perimeter=2(l+b)=2(7+9)=32units (ANS)

Regards.

Akash Kumar Dutta
10 years ago

Dear Vivek,
i am very sorry.
u see my first ans.
l=7 and b=9

so perimeter=32 cm.

i am very sorry if i have posted any wrong...those were mistyping.

Regards.